# Towards Context-Aware Domain Generalization: Understanding the Benefits and Limits of Marginal Transfer Learning

### v2

@@ 3.2 Learning Permutation-Invariant Representations
Analyzing set-structured data with neural networks has received much theoretical [Wagstaff et al., 2022, Bloem-Reddy and Teh, 2020, Murphy et al., 2018] and empirical [Zaheer et al., 2017, Lee et al., 2019, Zare and Van Nguyen, 2021] momentum in recent years. For instance, Zare and Van Nguyen [2021] build on the set transformer architecture [Lee et al., 2019] and augment the attentive encoder with the capability to learn dynamic templates for attention-based pooling. Differently, dan Guo et al. [2021] proposes to learn set-specific representations, along with global “prototypes”, using an optimal transport (OT) optimization criterion.
Notably, the methods above impose no probabilistic structure on the set-representations, since the latter are mainly used as deterministic features for downstream tasks. In the Bayesian literature, sets represent finitely exchangeable sequences, embodying the core probabilistic structure of most Bayesian models [Orbanz and Roy, 2014]. Typically, hierarchical (aka multi-level) Bayesian models are used to deal with Simpson’s paradox by organizing observations into clusters or levels [Wikle, 2003, Gelman et al., 2013], mirroring the notions of domain or environment in DG. Indeed, hierarchical Bayesian models have been successfully applied in many areas of science, but are typically constrained to linear or generalized linear models (GLMs), prioritizing interpretability over predictive performance.
The machine learning literature is rich with variational methods for learning invariant summary statistics from set-structured data [Edwards and Storkey, 2016, Garnelo et al., 2018, Kim et al., 2019, 2021, Zeng et al., 2022]. Crucially, none of the above methods pursues the concrete goal of our work, which is improving domain generalization performance through set-based environment representations. To achieve this goal, we employ variants of the DeepSet [Zaheer et al., 2017] and SetTransformer [Lee et al., 2019] architectures for our backbone set-encoder throughout all experiments. Interestingly, we observe that our approach is widely robust to the choice of set-encoder architectures in the problems considered.
Figure 2: Experiment 1. Left: Toy dataset that conforms to our theoretical criteria. A detailed description can be found in subsection E.1. Without environmental information, the marked input at $x=2.5$ could belong to either one of the domains numbered 1, 2, or 3. Marginal distributions of the five environments are shown on top. Right: Comparison of environment classification accuracy for different set sizes. Distances between environments refer to the distance between the means of the environments.
@@ 3.3 OOD Detection and Selective Classification
Detecting unusual inputs that deviate from the examples in the training set has been a long-standing problem of conceptual complexity in machine and statistical learning [Aggarwal and Yu, 2001, Yang et al., 2021, Shen et al., 2021, Han et al., 2022, Yang et al., 2022a]. Flagging out-of-distribution (OOD) instances involves identifying uncommon data points that might compromise the reliability of machine learning systems [Yang et al., 2021]. OOD detection is closely related to inference with a reject option (also termed selective classification) [Geifman and El-Yaniv, 2017, El-Yaniv et al., 2010], which allows classifiers to refrain from predicting ambiguous or novel conditions [Hendrickx et al., 2021]. The reject option has been extensively studied in statistical and machine learning [Hellman, 1970, Fumera and Roli, 2002, Grandvalet et al., 2008, Wegkamp and Yuan, 2012], with early work dating back to the 1950s [Chow, 1957, 1970, Hellman, 1970].
More recently, Müller et al. [2023] explored the utility of selective classification in DG settings. They investigated various post-hoc scores to define a “competence region” in feature space where a classifier is deemed competent. In this work, we consider a post-hoc score based on the $k$-nearest neighbours to the training set in feature space similar to Sun et al. [2022], which applies to both classification and regression settings. Unlike the approach taken in Müller et al. [2023], where the focus lies on features of individual instances, we consider the set summary provided by the set-encoder. Thus, we can identify novel environments even when singleton inputs lack sufficient information.
@@ 4 Experiments
In the following, we explore various aspects of the context-aware model across three different dimensions. First, we show on two datasets that a context-aware model achieves improved performance in the ID as well as the OOD setting compared to a baseline model when the conditions of a source component shift are met. Second, we show how novel environments can be detected to select between the most predictive (in the ID setting) and the most robust (in the OOD setting) model. We also show that novel environment detection can be utilized to avoid failure cases. Third, we demonstrate that the necessary criteria (see subsection 2.3) can be validated empirically, identifying cases where no benefits of the method can be expected. Experimental details can be found in the Appendix. Furthermore, in Appendix B, we conduct an additional experiment resembling Experiment 1, but with high-dimensional inputs
Figure 3: Experiment 1. Relative improvement of set-encoder (shown in I) approach versus baseline model (0 means no improvement is achieved) on a toy example. We also show I (OOD) on OOD data. II depicts the relative improvement of the environment-oracle model compared to the baseline model. III demonstrates the relative improvement in predicting the environment when using contextual information compared to the absence of it. Sampling variation arises from using different seeds to partition the ID data into training, test and validation set.
@@ 4.1 Evaluation Approach
To approximate Criterion 2.1, Criterion 2.2 and Criterion 2.3, we are required to train five distinct models (see Table 3 for an overview). We denote our composite model as $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ (see Figure 1) and the baseline model (having no access to the context) as $f^{\mathbf{Y}\mid\mathbf{X}}$. Based on these two models, we can compute the relative improvement achieved by our model relative to a baseline model via
$\mathcal{R}_{\text{I}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}.$ (2)
Here, $\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})$ denotes a performance measure for our model and similarly for the baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$, both measured on an unseen test set. $\mathcal{R}_{\text{I}}>0$ signifies an advantage attained by the context-aware approach and therefore the fulfillment of Criterion 2.1. In the regression setting, we consider the negative L2-Loss as the performance measure.
To validate Criterion 2.2, we train a contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ utilizing both the set input $\mathbf{S}^{(n)}$ and the target input $\mathbf{X}$ to predict the environment label $E$. Additionally, we train a baseline environment model $f^{E\mid\mathbf{X}}$ aimed at predicting $E$ solely from $\mathbf{X}$. We then compute the relative improvement $\mathcal{R}_{\text{II}}$ [^2] The exact definition for $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ can be found in Appendix D of the contextual environment predictor relative to the baseline environment predictor. $\mathcal{R}_{\text{II}}>0$ indicates that Criterion 2.2 is satisfied. In our experiments, we choose the set size $n$ such that we achieve approximately 100% accuracy for our contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ on ID data.
Similarly, we consider an environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ that aims to predict $\mathbf{Y}$ from the singleton input and the environment label $E$. We define the relative improvement $\mathcal{R}_{\text{III}}$ of the environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ compared to the baseline method $f^{\mathbf{Y}\mid\mathbf{X}}$. In this case, the relative improvement $\mathcal{R}_{\text{III}}$ is associated with Criterion 2.3.
@@ 4.2 Experiment 1: Toy Example
To set the stage, we consider a dataset shown in Figure 2 and explained in detail in subsection E.1. The dataset includes data from five different environments, defined by distinct Gaussian distributions. Each Gaussian deviates due to its location (i.e. mean vector). The dataset exemplifies Simpson’s paradox, wherein a naive fit of the data (without accounting for environmental factors) would yield a negatively sloped line (see Figure 10). Conversely, using environmental information for prediction enables to capture the true underlying model with positively sloped lines, counteracting the overall negative trend of the mean vectors.
Importantly, the dataset meets our necessary criteria: We cannot infer the origin environment from a single input alone, as indicated by the overlap between the marginal distributions of $X$, obtained by projecting the entire sample onto the $x$-axis in Figure 2. Thus, the setting aligns with Criterion 2.2, and corresponds to Criterion 2.3, since knowing the environment location $\boldsymbol{\mu}^{(j)}$ should improve prediction.
@@ Results
As a first check of Criterion 2.2, we evaluate whether a set input provides additional information about the environment compared to a singleton input. Figure 2 illustrates that additional set input improves the ability to distinguish between environments significantly and the more samples we include, the better the distinction. As expected, a decrease in the distance between environment marginal means necessitates more samples to differentiate between environments. Interestingly, the particular choice of architecture for the permutation-invariant network does not seem to play a significant role for predicting the environment label well, as demonstrated in Appendix I.
Next, we assess the predictive capabilities of the context-aware approach across all possible scenarios of “leave-one-environment-out”. This involves training on all environments except one and treating the excluded environment as a novel OOD scenario. Here, we consider linear models to ensure an optimal inductive bias for the problem (non-linear models achieve similar results, as shown in subsection E.3). We can see that Criterion 2.1, Criterion 2.2 and Criterion 2.3 are satisfied in Figure 3. Providing contextual information in the form of a set input increases the performance significantly compared to a baseline model in the ID as well as in the OOD setting (see I and I (OOD) in Figure 3). We also observe a slightly higher relative improvement when the environment label is directly provided (see II) compared to using the output of the set-encoder (see I). This aligns with our expectations, as the set input does not offer more information about the target value than the environment label itself. Note that for III we achieve less relative improvement since we consider the accuracy here and not the L2-Loss.
Finally, we visualize the predictions of the baseline approach and our set-encoder approach in Figure 4 for one trained model. Our model captures and utilizes the characteristics of each environment for prediction. In contrast, the baseline approach struggles to discern between environments due to the significant overlap between environments, resulting in an inability to deal with environmental differences. Note that we obtained the best results by considering a class of linear models that aligns with the data-generating process. However, we observe that extrapolation performance drops when the considered models are overly complex and lack a strong inductive bias (see subsection E.3).
4.3 Experiment 2: Colored MNIST
Setup
The ColoredMNIST dataset [Arjovsky et al., 2019] is an extension of the standard MNIST dataset, wherein the number of classes is reduced to two classes (all standard labels $<5$ are assigned to new label 0, and all labels $\geq 5$ are the new label 1). Furthermore, label noise is deliberately added, so only in 75% of all cases, the label can correctly be predicted from the shape. To make things more challenging, the image background can take two colors that are also associated with the image label. In the first environment, the association is 90% and in the second one 80%. Therefore, a baseline model would tend to utilize the background for prediction instead of the actual shape. However, in a third environment, the associations are reversed, so that a model based on the background color would achieve only 10% accuracy – worse than random.
This dataset implies a trade-off between predictive performance in ID domains versus robustness in OOD domains, as discussed in [Arjovsky et al., 2019, Zhang et al., 2023]. For instance, an invariant model that relies solely on an object’s shape would be robust to domain shift at the cost of diminished accuracy in the first two environments (75% instead of 80% or 90%). In contrast, a baseline model would achieve greater accuracy in the first domains (80% and 90%), but would fail dramatically in the third domain (only 10%).
[…]
@@ 4.3 Experiment 2: Colored MNIST
Bayes Optimal Classifier $\mathbf{85.0}$ $\mathbf{75.0}$
Table 1: Experiment 2. Mean and standard deviation of accuracy percentages across model types and domain settings over 5 runs. The features extracted by our model allow for improved OOD detection compared to the features of the baseline model. Thus, our model can perform a favorable selection between the baseline model in the ID setting and the invariant model in the OOD setting.
@@ Results
Here, we assume the invariant model to be given (see Appendix G for details), but it could also be obtained by invariant learning, e.g. Invariant Risk Minimization [Arjovsky et al., 2019]. With our novel environment detection approach (see subsection 2.5) we can get the best of both worlds, circumventing the inherent trade-off. When identifying the ID setting, we utilize the baseline model that achieves the highest predictiveness within the observed environments. In case we detect the OOD setting, we employ the invariant model. We compare this kind of model selection due to the features $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ inherent to our model versus the features extracted by the baseline model. The results can be found in Table 1. By utilizing the model selection based on the set-summary $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, we nearly recover the ID accuracy while maintaining identical performance to the invariant model on OOD data. Evidently, the novel environment detection only works with set summaries. A feature extracted from a single sample does not provide enough information to reliably detect distribution shifts, leading to difficulties in effectively selecting between baseline and invariant model, as demonstrated in Table 1.
@@ 4.4 Experiment 3: Violated Criteria
We consider the PACS dataset [Li et al., 2017], training our model on the Cartoons, Sketches, and Paintings environments, and assess its performance in the Art environment during testing. The dataset includes images with labels that we intend to predict. Regarding a second classification task, we delve into the OfficeHome dataset [Venkateswara et al., 2017]. In line with the PACS dataset, we approach the classification problem, training across three specific environments, and subsequently evaluating a novel one as an out-of-distribution (OOD) scenario.
@@
Figure 5: Experiment 3. Tell-tale examples where at least one of the necessary criteria is not satisfied and the context-aware approach cannot possibly yield benefits. For experimental details, see Appendix H.
@@ Results
When the criteria are not met, no benefits can be achieved, even in the ID setting. This has been proven in Theorem 2.1 and we demonstrate it here for two scenarios empirically (see Figure 5). We find that Criterion 2.2 is not satisfied on the PACS dataset: As depicted in 4(a), the contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ does not perform better compared to the baseline environment model $f^{E\mid\mathbf{X}}$. Remarkably, a single example is sufficient to infer the source environment, allowing for a 99.7% accuracy in predicting the correct environment from an individual sample (see Appendix H). Also Criterion 2.3 is not satisfied as 4(a) depicts. Since the criteria are not met, we do not achieve any benefit over the baseline model, neither in the ID nor in the OOD setting, as demonstrated in 4(a).
On the OfficeHome dataset we find that Criterion 2.2 is not satisfied, while Criterion 2.3 is. Results are depicted in 4(b). We observe that the set input offers benefits for predicting the data originating environment corresponding to Criterion 2.3. However, even when providing the target classifier with the environment label (environment-oracle model), we do not achieve an improvement over the baseline model suggesting that Criterion 2.2 is not satisfied. As expected and depicted in 4(b), our method does not yield benefits compared to the baseline model.
4.5 Experiment 4: Failure Case Detection
Setup
Besides unfulfilled criteria, another reason why a context-aware approach might fail to reap benefits is when the distribution shift requires extrapolation. This might be unattainable by the model, making the inclusion of a reject option beneficial. Using the BikeSharing dataset [Fanaee-T, 2013], we demonstrate that in cases where different seasons like summer or winter represent distinct environments, extrapolation might be necessary. For this dataset we consider the task of predicting the number of bikes rented across the day based on weather data. Here we explore the scenario where we train on all seasons except winter. Details about the dataset, pre-processing steps and other scenarios can be found in Appendix J.
Results
In Table 5 we demonstrate that the context-aware approach is slightly superior compared to the baseline model in the ID settings. However, both the baseline and the context-aware approach experience performance degradation in the novel winter environment. To detect the novel environment and, consequentially, potential failure cases, we compute the score as suggested in subsection 2.5 and evaluate how well it distinguishes between ID versus OOD environments. We designate an independent ID test set and use the environment excluded during training (here winter) as the OOD set for evaluation. The area under the ROC-curve (AUROC) in Table 5 demonstrates that the score based on the permutation-invariant embedding allows for a perfect detection of the novel environment, whereas the standard approach fails as expected.
Table 2: Experiment 4. Performance comparison between our model and the baseline for the winter domain in the BikeSharing dataset. We compare the performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC). Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case. We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets. Other domains can be found in Appendix J.

### v3

@@ 3.2 Learning Permutation-Invariant Representations
Analyzing set-structured data with neural networks has received much theoretical [31, 8, 32] and empirical [33, 34, 35] momentum in recent years. For instance, [35] build on the set transformer architecture [34] and augment the attentive encoder with the capability to learn dynamic templates for attention-based pooling. Differently, [36] proposes to learn set-specific representations, along with global “prototypes”, using an optimal transport (OT) optimization criterion.
A set-embedding can also be understood as a learned proxy variable for the confounder $E$. Generic proxy variables for confounding variables have been explored in the context of estimating the causal effect from $\mathbf{X}$ to $\mathbf{Y}$ in [37, 38]. While their work focuses on eliminating the effect of the confounding variable $E$, our objective is to leverage it for prediction purposes. Furthermore, they require $\mathbf{X}$ causing $\mathbf{Y}$ which does not conform to all prediction tasks. We do not require that $\mathbf{X}$ causes $\mathbf{Y}$ in our theoretical analysis and therefore include more scenarios (e.g., when $\mathbf{Y}$ is causing $\mathbf{X}$).
@@ 3.3 OOD Detection and Selective Classification
Detecting unusual inputs that deviate from the examples in the training set has been a long-standing problem of conceptual complexity in machine and statistical learning [39, 40, 41, 42, 43]. Flagging OOD instances involves identifying uncommon data points that might compromise the reliability of machine learning systems [40]. OOD detection is closely related to inference with a reject option (also termed selective classification) [44, 45], which allows classifiers to refrain from predicting ambiguous or novel conditions [46]. The reject option has been extensively studied in statistical and machine learning [47, 48, 49, 50], with early work dating back to the 1950s [51, 52, 47].
More recently, [9] explored selective classification in DG settings. They investigated various post-hoc scores to define a “competence region” in feature space where a classifier is deemed competent. In this work, we consider a post-hoc score based on the $k$-nearest neighbours to the training set in feature space similar to [53], which applies to both classification and regression settings. Unlike the approach taken in [9], where the focus lies on features of individual instances, we consider the set summary provided by the set-encoder. Thus, we can identify novel environments even when singleton inputs lack sufficient information.
@@ 4 Experiments
In the following, we explore various aspects of context-aware models. First, we show on two datasets that a context-aware model achieves improved performance in ID and OOD settings compared to a baseline model when the necessary conditions of a source component shift are met. Second, we show how novel environments can be detected to select between the most predictive (in the ID setting) and the most robust (in the OOD setting) model. We also show that novel environment detection can be utilized to avoid failure cases. Third, we demonstrate that the necessary criteria (see Section 2.3) can be validated empirically, identifying cases where no benefits of the method can be expected. Experimental details can be found in the Appendix and the source code is available at  [^1] https://github.com/LarsKue/context-aware-domain-generalization.
Figure 2: Experiment 1. Left: Toy dataset that conforms to our theoretical criteria. Without environmental information, the marked input at $x=2.5$ could belong to either one of the domains numbered 1, 2, or 3 as indicated by the marginal distributions shown on top. Right: Comparison of environment classification accuracy for a baseline model versus a mean-pooled set-encoder using different set sizes. Distances between environments refer to the distance between the means of the environments. Smaller distances produce stronger overlap of the marginals. A detailed description can be found in Section E.1.
@@ 4.1 Evaluation Approach
To approximate Criterion 2.1, Criterion 2.2, and Criterion 2.3, we train five models (Table 1). Our context-aware model $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ leverages context sets, while the baseline $f^{\mathbf{Y}\mid\mathbf{X}}$ ignores them. Their relative improvement is
$\mathcal{R}_{\text{I}}=\frac{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})},$ (2)
where $\mathcal{M}(\cdot)$ is a test performance metric (negative L2-loss for regression). $\mathcal{R}_{\text{I}}>0$ indicates that Criterion 2.1 holds. For Criterion 2.2, we compare the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ with its baseline $f^{E\mid\mathbf{X}}$:
$\mathcal{R}_{\text{II}}=\frac{\mathcal{M}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{M}(f^{E\mid\mathbf{X}})}{\mathcal{M}(f^{E\mid\mathbf{X}})}.$ (3)
We set $n$ such that $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ achieves nearly perfect ID accuracy; $\mathcal{R}_{\text{II}}>0$ supports Criterion 2.2. Finally, to test Criterion 2.3, we introduce the environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ and compute
$\mathcal{R}_{\text{III}}=\frac{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}.$ (4)
These relative improvement metrics serve as proxies for the theoretical criteria: when $\mathcal{M}$ is cross-entropy under optimal models, they align with mutual information measures. However, if $\mathcal{M}$ accuracy, then this proxy is not isomorphic to the criterion it approximates.
Model Symbol Description Purpose
Context-aware (ours) $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts $\mathbf{Y}$ from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Test Criterion 2.1
Baseline $f^{\mathbf{Y}\mid\mathbf{X}}$ Predicts $\mathbf{Y}$ from $\mathbf{X}$ Reference
Environment-oracle $f^{\mathbf{Y}\mid\mathbf{X},E}$ Predicts $\mathbf{Y}$ from $\mathbf{X}$ and $E$ Test Criterion 2.3
Contextual env. $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts $E$ from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Test Criterion 2.2
Baseline env. $f^{E\mid\mathbf{X}}$ Predicts $E$ from $\mathbf{X}$ Reference for Criterion 2.2
Table 1: Five models used to evaluate our approach and verify the theoretical criteria.
Figure 3: Experiment 1. Relative improvement of marginal transfer learning (shown in I) versus a baseline model (0 means no improvement is achieved) on a toy example. We also show I (OOD) on OOD data. II depicts the relative improvement of the environment-oracle model compared to the baseline model. III demonstrates the relative improvement in predicting the environment when using contextual information compared to the absence of it. Sampling variation arises from using different seeds to partition the ID data into training, test, and validation sets.
@@ 4.2 Experiment 1: Toy Example
To set the stage, we consider a dataset shown in Figure 2. The dataset includes data from five different environments, defined by distinct Gaussian distributions. Each Gaussian deviates due to its location (i.e. mean vector). The dataset exemplifies Simpson’s paradox, wherein fitting without accounting for environmental factors would yield a negatively sloped line. This trend reverses to multiple positively sloped lines when considering environmental factors (see Figure 11).
Importantly, the dataset meets our necessary criteria, since the environment cannot be inferred from a single input as indicated by the overlap of the marginal distributions in Figure 2. The mathematical details underlying this dataset are described in Section E.1.
@@ Results
As a first check of Criterion 2.2, we evaluate whether a set input provides additional information about the environment compared to a singleton input. Figure 2 illustrates that additional set input improves the ability to distinguish between environments significantly and the more samples we include, the better the distinction. As expected, a decrease in the distance between environment marginal means necessitates more samples to differentiate between environments.
Next, we assess the predictive capabilities of the context-aware approach across all possible scenarios of “leave-one-environment-out”. This involves training on all environments except one and treating the excluded environment as a novel OOD scenario. Here, we consider linear models to ensure an optimal inductive bias for the problem. We can see that Criterion 2.1, Criterion 2.2 and Criterion 2.3 are satisfied in Figure 3. Providing contextual information in the form of a set input increases the performance significantly compared to a baseline model in the ID as well as in the OOD setting (see I and I (OOD) in Figure 3). We also observe a slightly higher relative improvement when the environment label is directly provided (see II) compared to using the output of the set-encoder (see I). This aligns with our expectations, as the set input does not offer more information about the target value than the environment label itself. Note that for metric III, we achieve less relative improvement since we consider the accuracy and not the L2-Loss.
In Appendix E all scenarios where one environment is left out for testing can be found. Additionally, we present there similar results for non-linear models and also demonstrate that the specific choice of permutation-invariant network does not significantly impact the prediction of the environment label. Furthermore, in Appendix B, we conduct an additional experiment resembling Experiment 1, but with high-dimensional inputs and achieve similar results.
Accuracy $\%$] $\uparrow$
@@
Bayes Optimal $85.0$ $75.0$
Table 2: Experiment 2. Accuracy across model types and domain settings. Our context-aware model yields improved OOD detection compared to the baseline, allowing model selection at inference time. See Appendix K for more details.
4.3 Experiment 2: Colored MNIST
Setup
The ColoredMNIST dataset [54] is an extension of the standard MNIST dataset, wherein the number of classes is reduced to two classes (digits <5<5 and $\geq 5$). Furthermore, label noise is deliberately added, such that only in 75% of all cases, the label can be correctly predicted from the shape. To make things more challenging, the image background can take two colors that are also associated with the image label. In the first environment, the association is 90% and in the second one 80%. Therefore, a baseline model would tend to utilize the background for prediction instead of the actual shape. However, in a third environment, the associations are reversed, so that a model based on the background color would achieve only 10% accuracy (i.e., worse than random).
This dataset implies a trade-off between predictive performance in ID domains versus robustness in OOD domains, as discussed in [54, 30]. For instance, an invariant model that relies solely on an object’s shape would be robust to domain shift at the cost of lower accuracy in the first two environments (75% vs. 80% or 90%). In contrast, a baseline model would achieve greater accuracy in the first domains (80% and 90%), but would fail dramatically in the third domain (only 10%).
@@
Here, we assume the invariant model to be given, but it could also be obtained by invariant learning, e.g. Invariant Risk Minimization [54]. With our novel environment detection approach (see Section 2.5) we can get the best of both worlds, circumventing the inherent trade-off. When identifying the ID setting, we utilize the baseline model that achieves the highest predictiveness within the observed environments. In case we detect the OOD setting, we employ the invariant model. We compare this kind of model selection due to the features $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ inherent to our model versus the features extracted by the baseline model.
The results can be found in Table 2. By utilizing model selection based on the set-summary $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, we nearly recover the ID accuracy while maintaining identical performance to the invariant model on OOD data. Evidently, the novel environment detection only works with set summaries. A feature extracted from a single sample does not provide enough information to reliably detect distribution shifts, leading to difficulties in effectively selecting between baseline and invariant model, as demonstrated in Table 2. Details on this experiment can be found in Appendix G.
@@ 4.4 Experiment 3: Violated Criteria
To demonstrate the effects of criterion violation, we consider the PACS dataset [55], as well as the OfficeHome dataset [56], each with the Art environment chosen as the novel (OOD) domain.
Results
As expected, when the criteria are not met, context-aware models cannot achieve a benefit over the baseline (see Figure 4). Validating the criteria empirically, we find that Criterion 2.2 is not satisfied for PACS, as a single sample is sufficient to infer the source domain with near-perfect accuracy. Furthermore, Criterion 2.3 is not satisfied, as 4(a) depicts.
@@ Results
Figure 4: Experiment 3. Tell-tale examples where at least one of the necessary criteria is not satisfied and the context-aware approach cannot possibly yield benefits.
On the OfficeHome dataset we find that Criterion 2.2 is not satisfied, while Criterion 2.3 is. Results are depicted in 4(b). We observe that the set input offers benefits for predicting the source environment corresponding to Criterion 2.3. However, even when providing the target classifier with the environment label, we do not achieve a benefit, suggesting that Criterion 2.2 is not satisfied. For experimental details, see Appendix H.
4.5 Experiment 4: Failure Case Detection
Winter MSE $\downarrow$ AUROC $\%$] $\uparrow$
ID OOD
Baseline $2.21\pm 0.11$ $6.08\pm 0.13$ $58.2\pm 0.7$
Ours $\mathbf{2.09}\pm 0.12$ $\mathbf{5.7}\pm 0.4$ $\mathbf{100.0}\pm 0.0$
Table 3: Experiment 4. Inference performance (MSE) and novel environment detection (AUROC) comparison between our context-aware model and the baseline for the winter domain in the BikeSharing dataset. See Appendix K for more details.
Setup
Besides unfulfilled criteria, another reason why a context-aware approach might fail to reap benefits is when the distribution shift requires extrapolation. This might be unattainable by the model, making the inclusion of a reject option beneficial. Using the BikeSharing dataset [57], we demonstrate that in cases where different seasons like summer or winter represent distinct environments, extrapolation might be necessary. We consider the task of predicting the number of bikes rented across the day based on weather data. Here we explore the scenario where we train on all seasons except winter. Details about the dataset, pre-processing steps, and other testing scenarios can be found in Appendix J.
@@
In Table 3 we demonstrate that the context-aware approach is slightly superior compared to the baseline model in the ID settings. However, both the baseline and the context-aware approach experience performance degradation in the novel winter environment. To detect the novel environment and, consequentially, potential failure cases, we compute the score as suggested in Section 2.5 and evaluate how well it distinguishes between ID versus OOD environments. We designate an independent ID test set and use the environment excluded during training (here winter) as the OOD set for evaluation. The area under the ROC-curve (AUROC) in Table 3 demonstrates that the score based on the permutation-invariant embedding allows for perfect detection of the novel environment, whereas the standard approach fails as expected.